OpenStudy (anonymous):

Concrete can be purchased by the cubic yard. How much will it cost to pour a slab 11 feet by 11 feet by 3 inches for a patio if the concrete costs $63.00 per cubic yard?

OpenStudy (campbell_st):

convert all you measurements to yards 11ft = 11/3 yards 3 inches = 1/4 ft = 1/12 of a yard use these to find the volume in cubic yards V = length x width x height Cost = 63 x V